# Labrador: Exploring the limits of masked language modeling for laboratory data

### v1

@@ 1 Introduction
Code: https://github.com/DavidBellamy/labrador
In recent years, self-supervised pre-training of masked language models (MLMs) (see Appendix A for background) has demonstrated remarkable success across a wide range of machine learning problems and has led to significant downstream improvements across diverse tasks in natural language processing (Liu et al., 2019; Devlin et al., 2019; Raffel et al., 2020). There is considerable excitement surrounding the potential of large pre-trained MLMs to achieve similar success in medical applications. For instance, existing applications of MLMs in medicine have already yielded promising results in tasks related to medical text understanding (Lee et al., 2020; Alsentzer et al., 2019; Huang et al., 2019; Yang et al., 2019; Beltagy et al., 2019). Despite the success of self-supervised models in certain areas of biomedicine (Palepu & Beam, 2022; Beam et al., 2023; Singhal et al., 2023; Moor et al., 2023; Tiu et al., 2022; Ingraham et al., 2022; Watson et al., 2023; Shay et al., 2023), there has been no previous work on pre-trained models for laboratory measurements. Laboratory data is abundant, routinely collected, less biased compared to other types of data in electronic health records (EHRs) like billing codes (Beam et al., 2021), and directly measure a patient’s physiological state, offering a valuable opportunity for creating a medical foundation model.
@@
Contributions
  • We design a novel Transformer architecture, Labrador, that can be pre-trained on (integer, float) data in a truly continuous manner and show that it outperforms BERT.
  • We provide a convenient method to tokenize EHR datasets that include a mixture of categorical and continuous data using the empirical cumulative distribution function (eCDF) of each continuous feature.
  • We demonstrate that MLM pre-training on MIMIC-IV data does not transfer well to downstream outcome prediction tasks and discuss the implications for future pre-training work with lab data.
  • We provide key future directions for self-supervised pre-training on structured EHR data.
  • We open-source our outcome prediction evaluation datasets, model weights and code so that future work can be benchmarked against our results.
2 Background
@@
Rossi et al. (2019) learned embeddings of laboratory data using Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) by encoding each lab test result as the concatenation of a binary indicator of abnormality and the test’s LOINC code [^1] https://loinc.org/get-started/what-loinc-is/. The ablations showed improved mortality prediction when the binary indicator was included in the embeddings despite this being only a binary representation of the underlying continuous lab test result.
@@
Figure 1: Labrador model architecture.
@@ 3 Methods
In this section, we describe the Labrador and BERT model architectures and pre-training. For further experimental details, see Appendix C.
@@ 3.1 Labrador architecture
Labrador consists of two embedding modules (categorical and continuous), N transformer blocks, and two prediction heads (categorical and continuous). The categorical embedding layer uses the standard integer lookup method, whereas the continuous embedding layer is defined in the following pseudocode:
def ContinuousEmbeddingLayer(lab_values, token_embeddings):
    x = Dense_pw(lab_values, out_size=d_model, activation=‘linear’)
    x = x + token_embeddings
    x = Dense_pw(x, out_size=d_model, activation=‘relu’)
    x = LayerNorm(x)
    return x
@@
The categorical prediction head consists of a ReLU-activated position-wise dense layer followed by a softmax layer over the 529 token vocabulary. The pseudocode for the continuous prediction head is below:
def ContinuousPredictionHead(final_layer_embeddings, probs):
    x = concat([final_layer_embeddings, probs])
    x = Dense_pw(x, out_size=embed_dim + probs_dim, activation=‘relu’)
    x = Dense_pw(x, out_size=1, activation=‘sigmoid’)
    return x
Where probs is the categorical head’s probability distribution over the lab code vocabulary. This supplies the continuous prediction head with each token’s predicted probability when attempting to impute the masked lab value. The output of this layer is simply a float on the interval $[0,1]$, since all input values are standardized to this interval using the empirical cumulative distribution function (see Section 3.3). In our evaluations, we use the following model configuration resulting in 196,642,645 parameters:
  • embedding_dim = 1024
  • N = 10
  • attention_heads = 4
  • feedforward_dim = 1024
@@ 3.2 BERT architecture
The model configuration (HuggingFace’s BERTConfig) that is used in our experiments is:
  • vocab_size: 4251
  • hidden_size: 1024
  • num_hidden_layers: 10
  • num_attention_heads: 4
  • intermediate_size: 1024
  • hidden_act: ReLU
@@
This BERT instance has 68,522,139 parameters, which is less than Labrador due to their difference in key_dim size. To ensure that this difference in parameter count does not impact our conclusions, we performed all experiments with a parameter-matched BERT that uses the same key_dim as Labrador (Appendix B). The results from these experiments are indistinguishable from the results in Section 4.
3.3 Preprocessing of the pre-training lab data
@@ 3.4 Transformer pre-training
Both Labrador and BERT were pre-trained using a single 40GB A100 GPU for 500,000 steps (80 hours) and 1.5M steps (237 hours), respectively. Our stopping criterion for each model was to observe that the validation loss was sufficiently converged. We also wanted to ensure that BERT’s degree of convergence was greater than or equal to Labrador’s in order to avoid falsely representing the capabilities of BERT in downstream evaluations. The following hyperparameters were used for both models:
  • Embedding dimension: 1024
  • Dropout rate: 0.1
  • Learning rate: 1e-5
  • Batch size: 256
@@ 4.1.2 Embedding space visualization
Despite the fact that BERT and Labrador have comparable pre-training perplexity, the structure of their embedding space differs greatly. We performed dimensionality reduction on BERT and Labrador’s embeddings for the labs in the test split using the UMAP algorithm (McInnes et al., 2018). Section C.1 describes the details of the experiment. Figure 2 presents the two-dimensional structure of the 70 most frequent lab tests, which are colored according to the panel of labs that they are most often ordered with. Appendix F defines each lab panel.
@@
It is important to note that while low-dimensional representations of embeddings can be helpful in providing high-level, qualitative intuitions for how a model behaves, they can be misleading and unfaithful to the original, high-dimensional space (Chari & Pachter, 2023). We provide these visualizations to provide some insight into the different ways Labrador and BERT learn to represent the input space, but caution should be used when attempting to make fine-grained, quantitative conclusions.
Figure 2: UMAP Visualization of Labrador and BERT Embeddings. A. A global view of the embedding space structure for Labrador (left) and BERT (right). The 70 most frequently ordered lab tests are shown and colored according to the panel of tests they are typically ordered with (see Appendix F for all panel definitions). All labs on this figure are from the test split and were not seen during pre-training. B. Visualization of embeddings for four routinely collected laboratory measurements colored by lab value and scaled to the interval [0, 1]. Labrador appears to encode the measured lab value in a much more natural way with a smooth gradient for lab value compared to BERT.
@@ 4.1.3 Imputation
We also evaluated Labrador and BERT for their ability to impute missing lab values in the test split of their pre-training data. To do so, we mask a randomly selected lab value from each bag of labs in the test split and run each model’s forward pass to obtain its prediction for the masked value (see Section C.2 for more details). In panel A of Figure 3, we see that the predictions from both pre-trained models achieve a Pearson correlation $r^{2}>0.8$. As an ablation, we also evaluated Labrador and BERT with randomly initialized parameters to assess the effect of pre-training on the imputations. We see that the strong correlation between the imputed and true lab values stems entirely from pre-training and is not an artifact of the model architectures or the data.
@@ 4.2.1 Fine-tuning experiments and outcome prediction
We compare Labrador and BERT’s performance to a set of baseline models in four distinct outcome scenarios where predictions were made on the basis of laboratory measurements: ICU mortality due to sepsis, cancer diagnosis, COVID-19 diagnosis, and the prediction of alcohol consumption. Baseline methods are defined in Section C.5 and pre-processing for each evaluation dataset is described in Section C.3.
For each of these experiments, we add a prediction head to Labrador or BERT consisting of multiple fully-connected layers followed by extensive hyperparameter tuning for this prediction head on each outcome dataset and with each base model. Section C.4 describes this fine-tuning procedure in detail. Table 1 shows the performance of Labrador and BERT compared to simple logistic (or linear) regression, Random Forest and XGBoost.
XGBoost narrowly outperforms all other methods on sepsis mortality prediction, cancer diagnosis and COVID-19 diagnosis. Labrador performs best on alcohol consumption prediction, which is a regression problem (as opposed to the other 3 classification tasks) and is also the smallest outcome dataset. Additionally, Labrador outperforms BERT across all four evaluations, suggesting there is value in a model that explicitly accounts for the continuous nature of the data. For a precise definition of each evaluation task see Section C.3.
@@
Sepsis
mortality ($\downarrow$)
Cancer
diagnosis ($\downarrow$)
COVID-19
diagnosis ($\downarrow$)
Alcohol
consumption ($\downarrow$)
LR1
0.410 (0.398, 0.424)
1.177 (1.100, 1.272)
0.462 (0.434, 0.529)
9.73 (4.83, 13.93)
RF
0.387 (0.372, 0.401)
0.942 (0.902, 0.987)
0.431 (0.404, 0.487)
9.65 (4.62, 13.35)
XGB
0.387 (0.376, 0.399)
0.918 (0.847, 0.989)
0.419 (0.371, 0.504)
10.84 (5.54, 15.39)
BERT
0.406 (0.368, 0.431)
1.01 (0.859, 1.127)
0.441 (0.391, 0.513)
8.34 (5.80, 12.71)
Labrador
0.400 (0.371, 0.420)
0.978 (0.889, 1.163)
0.425 (0.369, 0.487)
7.11 (5.33, 9.15)
@@
    LR = logistic regression for all evaluations but alcohol consumption, in which case LR = linear regression.
@@ 4.2.2 Ablation study
Pre-trained
Random
weights (%)11footnotemark: 1
Pre-trained
Random
weights (%)11footnotemark: 1
@@ 5 Discussion
  1. 1.
    Laboratory data lack the complex correlation structure that has made MLM successful on other data modalities.
@@
    If lab values were statistically independent then the MLM objective would be unsuitable and the pre-training task would be impossible. Likewise, if there is a trivial correlation structure (e.g. some pairs of labs are perfectly correlated) then we would not expect an MLM objective to outperform simple models. However, we believe that there is a meaningful and complex correlation structure amongst the various lab measurements, as evidenced in Figure 8.
  2. 2.
    The masking rate of the MLM pre-training objective is too low.
    Perhaps Labrador and BERT are able to optimize the pre-training objective without learning sufficiently useful representations of labs and patient physiology to deliver significant downstream value on outcome prediction. During our pre-training, only one lab test is masked per bag, which results in an average mask rate of approximately 5% per bag. However, half of the bags in the training data contain greater than 20 elements and it is possible that the mask rate is effectively too low. Unfortunately, increasing the mask rate may be impossible because each bag of labs is permutation invariant. As a result, if 2 or more elements are masked within a bag, Labrador and BERT’s predictions for each mask token will be identical. There is no way for the model to distinguish one mask token from the next in a permutation invariant setting. This is in contrast to language modeling, where there are positional embeddings for each index in the input sequence, enabling the model to differentiate mask tokens based on their relative positions in the input.
@@
    Positional embeddings are the most common approach to breaking the symmetry in this architecture, however we are not currently aware of a suitable variable that could be encoded using positional embeddings. For example, the index in the input cannot be used to encode positional information since the relative ordering of multiple mask tokens is arbitrary. Another option is to use the timestamp when a lab was measured to encode positional information and to turn a bag of labs into a time series of labs. However, labs are ordered in sets where every lab in the order set is recorded at the exact same moment in time in a patient’s medical record. Even if a bag of labs includes tests that span multiple order sets, there would still be a symmetry issue within each order set. We believe that this is a general challenge for any MLM of electronic health record data, since measurements are often made in order sets.
  3. 3.
    The pre-training and downstream evaluations are a poor match.
@@
    The set of features for the cancer diagnosis evaluation is the least similar to the pre-training vocabulary and is also the evaluation that both Labrador and BERT perform the worst on. However, the sepsis mortality prediction task uses lab data from the same ICU data distribution as in pre-training and yet it is not the evaluation where pre-training delivers the most value, which would be COVID-19 diagnosis. Therefore, we do not believe that a mismatch between pre-training and downstream data distributions is the primary factor explaining the failure in transfer learning.
  4. 4.
    Laboratory data only partially capture patient state.
@@
    Laboratory data, unlike text, do not provide a comprehensive view of a patient's condition, necessitating the incorporation of additional patient data categories. Thus, it could be that even simple models are close to the Bayes error rate given the information captured by lab measurements. We propose that future self-supervised pre-training should attempt joint learning from multiple EHR data categories, like diagnostic codes, procedure codes, and medication prescriptions. This could lead to a multimodal model predicting next tokens across structured EHR data, medical images, and clinical text, albeit presenting the challenge of organizing this data into coherent input sequences.
  5. 5.
    Insufficient data scale.
@@
    Many of the capabilities of today’s frontier deep learning systems, such as GPT-4 (OpenAI, 2023), have emerged unexpectedly as a function of the scale of the pre-training data. We pre-trained Labrador and BERT on approximately 100 million lab tests but this only corresponds to 4.5 million input samples (i.e. bags of labs). In comparison, GPT-3 (Brown et al., 2020) was trained on a corpus of 300 billion tokens where each input sequence contained 2048 tokens (the model’s maximum context-window length), or about 145 million input sequences. This is nearly two orders of magnitude larger than our pre-training dataset in terms of input sequences and over 3 orders of magnitude larger in token count. We believe that this is at least 1 order of magnitude larger than the largest available (private or public) lab dataset today. Therefore, we believe that another important future direction for the creation of foundation models for medicine is the harmonization of datasets across many sources, which requires large organized coding efforts to unify the data into a compatible vocabulary. This also requires open access to such data sources, as MIMIC has provided. Importantly, if we take into account the Chinchilla scaling laws from Hoffmann et al. (2022), it is plausible that scaling model size will not yield significant benefits in downstream performance given the relative scarcity of laboratory data compared to text.

### v2

@@ Abstract
keywords:
Transformer, EHR, lab data
Data and Code Availability
The MIMIC-IV database can be accessed via PhysioNet after completing a short online training. The lab data that was used to pre-train Labrador and BERT come from the labevents table within MIMIC-IV and our pre-processing code can be found in our codebase. The COVID-19, cancer diagnosis and alcohol consumption fine-tuning datasets are public and therefore we provide our pre-processed versions of these datasets directly in our codebase.
The codebase also contains all of the code necessary to replicate the pre-training data pre-processing, model pre-training, fine-tuning and other evaluations. We cannot directly share the sepsis dataset due to MIMIC-IV’s data sharing policy, however with access to MIMIC-IV the steps described in Section F.3.2 recreate our evaluation dataset. Finally, the weights for the pre-trained Labrador and BERT models can be downloaded from our HuggingFace repository here.
@@ 1 Introduction
In recent years, self-supervised pre-training of masked language models (MLMs) (see Appendix A for background) has demonstrated remarkable success across a wide range of machine learning problems and has led to significant downstream improvements across diverse tasks in natural language processing (Liu et al., 2019; Devlin et al., 2019; Raffel et al., 2020). There is considerable excitement surrounding the potential of large pre-trained MLMs to achieve similar success in medical applications. For instance, existing applications of MLMs in medicine have already yielded promising results in tasks related to medical text understanding (Lee et al., 2020; Alsentzer et al., 2019; Huang et al., 2019; Yang et al., 2019; Beltagy et al., 2019). Despite the success of self-supervised models in certain areas of biomedicine (Palepu and Beam, 2022; Beam et al., 2023; Singhal et al., 2023; Moor et al., 2023; Tiu et al., 2022; Ingraham et al., 2022; Watson et al., 2023; Shay et al., 2023), there has been no previous work on pre-trained models for laboratory measurements. Laboratory data is abundant, routinely collected, less biased compared to other types of data in electronic health records (EHRs) like billing codes (Beam et al., 2021), and directly measure a patient’s physiological state, offering a valuable opportunity for creating a medical foundation model.
@@
Figure 1: Labrador model architecture.
2 Related Work
@@
Rossi et al. (2019) learned embeddings of laboratory data using Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) by encoding each lab test result as the concatenation of an octal indicator of abnormality and the test’s LOINC code [^1] https://loinc.org/get-started/what-loinc-is/. The ablations showed improved mortality prediction when the octal indicator was included in the embeddings despite this being only a discrete representation of the underlying continuous lab test result.
@@
Recent work by Zhu et al. (2024) evaluated 14 language models alongside conventional clinical predictive models and found that even with well-designed prompts, the best LLMs’ zero-shot performance does not consistently beat traditional machine learning models when there are 100 or more examples to train on.
@@ 3 Methods
In this section, we describe the Labrador and BERT model architectures and pre-training. For further experimental details, see Appendix F.
BERT is primarily distinguished from other Transformers by its learning objective rather than its model architecture. BERT’s architecture is classified as encoder-only and consists of an embedding layer, N transformer blocks with the canonical layers (MultiHeadAttention, Dropout, LayerNorm, Feedforward, Dropout, LayerNorm), and a prediction head.
The primary contribution of the original work was the approach of masked language modeling (MLM). The key advantage of MLM is that it does not rely on the input data having a natural sequential ordering, unlike autoregressive language modeling. The lab data in our study does not have a natural sequential ordering so we had to use an MLM-based approach. We considered ways to add positional embeddings to the lab data such that they would have a sequential ordering (even if only artificially), but we concluded that it is impossible to find a natural ordering because batches of lab tests are requested at the same moment in time by physicians. So there is no way to order labs sequentially besides alphabetically or some other arbitrary order. We did not think that adding such arbitrary ordering would be helpful.
Besides encoder-only transformer architectures, GPT-style models have decoder-only architectures and older seq2seq transformer models (e.g. t5) have encoder-decoder architectures. Both decoder-only and encoder-decoder architectures are useful for generating sequences like text. In other words, these architectures were designed for seq2seq tasks. But since lab data do not have a sequential ordering, we cannot pose it as a seq2seq task so these transformer architectures are not applicable to our setting. This is why we chose a BERT-like (i.e. encoder-only) MLM-based approach in this work.
@@ 3.1 Labrador architecture
Labrador consists of two embedding modules (categorical and continuous), N transformer blocks, and two prediction heads (categorical and continuous). The categorical embedding layer uses the standard integer lookup method, whereas the continuous embedding layer involves a position-wise linear projection. A pseudo-code implementation is in Section B of the Appendix.
@@
The categorical prediction head consists of a ReLU-activated position-wise dense layer followed by a softmax layer over the 529 token vocabulary. A pseudo-code implementation of the continuous prediction head is also in Section B of the Appendix. The output of this layer is simply a float on the interval $[0,1]$, since all input values are standardized to this interval using the empirical cumulative distribution function (see Section 3.3). In our evaluations, we use 10 transformer blocks, an embedding dimension of 1024, 4 attention heads, and a feedforward dimension of 1024, resulting in 196,642,645 parameters.
@@ 3.2 BERT architecture
The model configuration (HuggingFace’s BERTConfig) that is used in our experiments has a vocabulary size of 4251, a hidden size of 1024, 10 hidden layers, 4 attention heads and an intermediate (i.e. feedforward) size of 1024 with a ReLU activation.
@@
This BERT instance has 68,522,139 parameters, which is less than Labrador due to their difference in key_dim size. To ensure that this difference in parameter count does not impact our conclusions, we performed all experiments with a parameter-matched BERT that uses the same key_dim as Labrador (Appendix E). The results from these experiments are indistinguishable from the results in Section 4.
3.3 Preprocessing of pre-training lab data
@@ 3.4 Transformer pre-training
Both Labrador and BERT were pre-trained using a single 40GB A100 GPU for 500,000 steps (80 hours) and 1.5M steps (237 hours), respectively. Our stopping criterion for each model was to observe that the validation loss was sufficiently converged. We also wanted to ensure that BERT’s degree of convergence was greater than or equal to Labrador’s in order to avoid falsely representing the capabilities of BERT in downstream evaluations. Both models were pre-trained with a dropout rate of 0.1, learning rate of 1e-5, a batch size of 256, and an embedding size of 1024.
@@ 4.1.2 Embedding space visualization
Despite the fact that BERT and Labrador have comparable pre-training perplexity, the structure of their embedding space differs greatly. We performed dimensionality reduction on BERT and Labrador’s embeddings for the labs in the test split using the UMAP algorithm (McInnes et al., 2018). Section F.1 describes the details of the experiment. Figure 2 presents the two-dimensional structure of the 70 most frequent lab tests, which are colored according to the panel of labs that they are most often ordered with. Appendix I defines each lab panel.
@@
It is important to note that while low-dimensional representations of embeddings can be helpful in providing high-level, qualitative intuitions for how a model behaves, they can be misleading and unfaithful to the original, high-dimensional space (Chari and Pachter, 2023). We provide these visualizations to provide some insight into the different ways Labrador and BERT learn to represent the input space, but caution should be used when attempting to make fine-grained, quantitative conclusions.
@@ 4.1.3 Imputation
We also evaluated Labrador and BERT for their ability to impute missing lab values in the test split of their pre-training data. To do so, we mask a randomly selected lab value from each bag of labs in the test split and run each model’s forward pass to obtain its prediction for the masked value (see Section F.2 for more details). In panel A of Figure 3, we see that the predictions from both pre-trained models achieve a Pearson correlation $r^{2}>0.8$. As an ablation, we also evaluated Labrador and BERT with randomly initialized parameters to assess the effect of pre-training on the imputations. We see that the strong correlation between the imputed and true lab values stems entirely from pre-training and is not an artifact of the model architectures or the data.
@@ 4.2.1 Fine-tuning experiments and outcome prediction
We compare Labrador and BERT’s performance to a set of baseline models in four distinct outcome scenarios where predictions were made on the basis of laboratory measurements: ICU mortality due to sepsis, cancer diagnosis, COVID-19 diagnosis, and the prediction of alcohol consumption. Baseline methods are defined in Section F.5 and pre-processing for each evaluation dataset is described in Section F.3.
For each of these experiments, we add a prediction head to Labrador or BERT consisting of multiple fully-connected layers followed by extensive hyperparameter tuning for this prediction head on each outcome dataset and with each base model. Section F.4 describes this fine-tuning procedure in detail. Table 1 shows the performance of Labrador and BERT compared to simple logistic (or linear) regression, Random Forest and XGBoost.
XGBoost narrowly outperforms all other methods on sepsis mortality prediction, cancer diagnosis and COVID-19 diagnosis. Labrador performs best on alcohol consumption prediction, which is a regression problem (as opposed to the other 3 classification tasks) and is also the smallest outcome dataset. Additionally, Labrador outperforms BERT across all four evaluations, suggesting there is value in a model that explicitly accounts for the continuous nature of the data. For a precise definition of each evaluation task see Section F.3.
@@
Sepsis mortality ($\downarrow$) Cancer diagnosis ($\downarrow$) COVID-19 diagnosis ($\downarrow$) Alcohol consumption ($\downarrow$)
LR1 0.410 (0.398, 0.424) 1.177 (1.100, 1.272) 0.462 (0.434, 0.529) 9.73 (4.83, 13.93)
RF 0.387 (0.372, 0.401) 0.942 (0.902, 0.987) 0.431 (0.404, 0.487) 9.65 (4.62, 13.35)
XGB 0.387 (0.376, 0.399) 0.918 (0.847, 0.989) 0.419 (0.371, 0.504) 10.84 (5.54, 15.39)
BERT 0.406 (0.368, 0.431) 1.01 (0.859, 1.127) 0.441 (0.391, 0.513) 8.34 (5.80, 12.71)
Labrador 0.400 (0.371, 0.420) 0.978 (0.889, 1.163) 0.425 (0.369, 0.487) 7.11 (5.33, 9.15)
@@
    LR = logistic regression for all evaluations but alcohol consumption, where LR = linear regression.
@@ 4.2.2 Ablation study
Pre-trained Random weights (%)11footnotemark: 1 Pre-trained Random weights (%)11footnotemark: 1
@@ 5 Discussion
Lab data lack correlation structure. If lab values were statistically independent then the pre-training task would be impossible. Likewise, if there is a trivial correlation structure (e.g. some pairs of labs are perfectly correlated) then we would not expect an MLM objective to outperform simple models. However, we believe that there is a meaningful and complex correlation structure amongst the lab measurements, as evidenced in Figure 8.
MLM masking rate is too low. Labrador and BERT may satisfy the pre-training objective without learning enough about labs and patients to be adept at outcome prediction. During our pre-training, the average mask rate is just 5% per bag. Unfortunately, increasing the mask rate may be impossible because each bag of labs is permutation invariant. As a result, if multiple elements are masked within a bag, Labrador and BERT’s predictions for each mask token will be identical. There is no way for the model to distinguish one mask token from the next in a permutation invariant setting.
@@
Positional embeddings are typically used to break this symmetry, but we are not aware of a variable to encode using via positional embeddings. For example, the index in the input cannot be used since the relative ordering of multiple mask tokens is arbitrary. The timestamp when a lab was measured could be used but labs are ordered in sets at the exact same moment in time in a patient’s medical record. Even if a bag of labs includes tests that span multiple order sets, there would still be a symmetry issue within each order set. We believe that this is a general challenge for any MLM of electronic health record data, since measurements are often made in order sets.
@@
Pre-training and downstream evaluations are a poor match. The set of features used for cancer diagnosis is the least similar to the pre-training vocabulary and is also the evaluation that both Labrador and BERT perform the worst on. However, sepsis mortality prediction uses lab data from the same ICU data distribution as in pre-training and yet it is not the evaluation where pre-training delivers the most value, which would be COVID-19 diagnosis. Therefore, we do not believe that a mismatch between pre-training and downstream data distributions is the primary factor explaining the failure in transfer learning.
@@
Lab data only partially capture patient state. Lab data, unlike text, do not provide a comprehensive view of a patient's condition, necessitating the incorporation of additional patient data categories. Thus, it could be that even simple models are close to the Bayes error rate given the information captured by lab measurements. We propose that future work should jointly learn from multiple EHR data categories, like diagnostic codes, procedure codes, and medication prescriptions. This could lead to a multimodal model predicting next tokens across structured EHR data, medical images, and clinical text.
@@
Insufficient data scale. The capabilities of frontier LLM’s have emerged as a function of the scale of the pre-training data. We pre-trained Labrador and BERT on approximately 100 million lab tests but this only corresponds to 4.5 million input samples. In comparison, GPT-3 (Brown et al., 2020) was trained on a corpus of 300 billion tokens where each input sequence contained 2048 tokens or about 145 million input sequences. This is nearly two orders of magnitude larger than our pre-training dataset in terms of input sequences and over 3 orders of magnitude larger in token count. We believe that this is at least 1 order of magnitude larger than the largest available lab dataset today. Therefore, we encourage others to contribute to the harmonization of datasets across many sources, which requires large organized coding efforts to unify the data into a compatible vocabulary. This also requires open access to such data sources, as MIMIC has provided. Importantly, if we take into account the Chinchilla scaling laws from Hoffmann et al. (2022), it is plausible that scaling model size will not yield significant benefits in downstream performance given the relative scarcity of lab data compared to text.